# Enhancing Automated Interpretability with Output-Centric Feature Descriptions

### v1

@@ Acknowledgements
We thank the Transluce team, specifically Dami Choi, for sharing their neuron description pipeline data, as well as Johnny Lin from Neuronpedia for sharing their descriptions and model activations data. This work was supported in part by the Gemma 2 Academic Research Program at Google, the Edmond J. Safra Center for Bioinformatics at Tel Aviv University, and the Israel Science Foundation grant 1083/24. Figures 1 and 2 use images from www.freepik.com.
@@ B.2 Description Generation
We tried different prompts, but didn’t observe significant improvement. These include both generic prompts to be used for all subject models (Figures 13 and 14), and more fine-tuned prompts based on vocabulary projection demonstrations for each subject model (see the fine-tuned based prompt in Figure 11, for which we concatenate few-shot examples for each model as seen in Figure 9).
@@ B.3 Recreating Neuronpedia Descriptions using MaxAct
Llama-3.1 MLP SAE GPT2 Res. Mid. SAE GPT2 Res. SAE GPT2 MLP SAE
Input Output Input Output Input Output Input Output
MaxAct 56.4 $\pm$ 2.9 49.6 $\pm$ 2.9 43.8 $\pm$ 4.4 45.5 $\pm$ 4.5 42.4 $\pm$ 4.5 43.9 $\pm$ 4.5 41.3 $\pm$ 4.4 32.6 $\pm$ 4.2
VocabProj 20.2 $\pm$ 2.3 48.2 $\pm$ 2.9    23 $\pm$ 3.8 46.1 $\pm$ 4.5 19.2 $\pm$ 3.5 43.7 $\pm$ 4.5   6.9 $\pm$ 2.3 37.4 $\pm$ 4.4
TokenChange 25.4 $\pm$ 2.5 53.1 $\pm$ 2.9 25.3 $\pm$ 3.9 42.8 $\pm$ 4.4 22.4 $\pm$ 3.8 42.6 $\pm$ 4.5   6.1 $\pm$ 2.1 34.9 $\pm$ 4.3
EnsembleR (MA+VP) 62.1 $\pm$ 2.8 45.8 $\pm$ 2.9 56.8 $\pm$ 4.4 45.7 $\pm$ 4.5 56.8 $\pm$ 4.5 46.4 $\pm$ 4.5 50.9 $\pm$ 4.5 36.1 $\pm$ 4.3
EnsembleR (MA+TC) 65.8 $\pm$ 2.8 48.9 $\pm$ 2.9 56.2 $\pm$ 4.4 46.6 $\pm$ 4.5 58.9 $\pm$ 4.4 49.4 $\pm$ 4.5 50.5 $\pm$ 4.5    40 $\pm$ 4.4
EnsembleR (VP+TC) 22.6 $\pm$ 2.4 50.7 $\pm$ 2.9 28.6 $\pm$ 4.1 44.3 $\pm$ 4.5 25.3 $\pm$ 3.9 44.3 $\pm$ 4.5   7.1 $\pm$ 2.3 43.6 $\pm$ 4.5
EnsembleR (All) 62.7 $\pm$ 2.8 51.6 $\pm$ 2.9    61 $\pm$ 4.4 47.2 $\pm$ 4.5   59.1 $\pm$ 4.4 48.7 $\pm$ 4.5 50.1 $\pm$ 4.5 36.9 $\pm$ 4.3
EnsembleC (All) 39.1 $\pm$ 2.8 55.5 $\pm$ 2.9 40.1 $\pm$ 4.4 49.5 $\pm$ 4.5   39.2 $\pm$ 4.4 46.4 $\pm$ 4.5 25.2 $\pm$ 3.9 37.7 $\pm$ 4.4
@@ Appendix C Additional Evaluation Results
Appendix D Additional Details and Examples for Dead Feature Analysis
D.1 Generating Candidate Prompts
@@
D.2 Dead Feature Revival Example
@@
Appendix E Additional Examples for Qualitative Analysis
@@
Appendix F Resources and Packages
In our experiments, we used models, data and code from the following packages: transformers Wolf (2019), datasets Lhoest et al. (2021), TransformerLens Nanda and Bloom (2022) and SAELens Joseph Bloom and Chanin (2024). All of the experiments were conducted using a single A100 80GB or H100 80GB GPU.
@@ Appendix F Resources and Packages
Figure 12: An example of steered text set for the output-based metric.

### v2

@@ Acknowledgements
We thank the Transluce team, specifically Dami Choi, for sharing their neuron description pipeline data, as well as Johnny Lin from Neuronpedia for sharing their descriptions and model activations data. This work was supported in part by the Gemma 2 Academic Research Program at Google, the Edmond J. Safra Center for Bioinformatics at Tel Aviv University, a grant from Open Philanthropy, and the Israel Science Foundation grant 1083/24. Figures 1 and 2 use images from www.freepik.com.
@@ Appendix A Additional Details on Feature Description Evaluations
Method Estimated FLOPs
Gemma-2 2B Gemma-2 9B Gemma-2 27B
MaxAct $3.9\cdot 10^{16}$ $1.5\cdot 10^{17}$    $5\cdot 10^{17}$
VocabProj $2.8\cdot 10^{14}$ $1.1\cdot 10^{15}$    $2\cdot 10^{15}$
TokenChange $9.9\cdot 10^{13}$ $4.1\cdot 10^{14}$ $1.3\cdot 10^{15}$
Table 4: Estimated FLOPs for generating descriptions for all MLP features for models of different sizes, on a sample of 25k sequences of 128 tokens each, as done by Neuronpedia.
@@
The clamping values for $m$ were derived by fixing two target KL-divergence values, $0.25$ and $0.5$, providing two positive and two negative clamping values for $m$. These values, along with sequence length, balance generating text with sufficient feature effect, and producing long or degenerate text that is difficult to evaluate. To confirm this, we ran the output-based evaluation using target KL-divergence values of $0.5$ and $1$ on 800 features from Gemma-2 2B, obtaining similar results. However, the generated text became more degenerate (see Figure 16 for an example text). Therefore, we decided to retain the original target KL-divergence values, as higher values resulted in text that did not reflect probable model behavior.
@@ B.1 Variants of VocabProj
Figure 5: Human evaluation results for 100 features across three methods, for input- and output-faithfulness.
Figure 6: Instructions provided to human annotators for the evaluation of feature descriptions. These were accompanied with a few example annotations.
@@ B.2 Description Generation
We tried different prompts, but didn’t observe significant improvement. These include both generic prompts to be used for all subject models (Figures 15 and 17), and more fine-tuned prompts based on vocabulary projection demonstrations for each subject model (see the fine-tuned based prompt in Figure 13, for which we concatenate few-shot examples for each model as seen in Figure 11).
@@ B.3 Recreating Neuronpedia Descriptions using MaxAct
Llama-3.1 MLP SAE GPT2 Res. SAE GPT2 MLP SAE
Input Output Input Output Input Output
MaxAct 56.4 $\pm$ 2.9 49.6 $\pm$ 2.9 44.4 $\pm$ 2.3 44.1 $\pm$ 2.3 39.7 $\pm$ 2.9    34 $\pm$ 2.8
VocabProj 20.2 $\pm$ 2.3 48.2 $\pm$ 2.9 23.7 $\pm$ 2 42.8 $\pm$ 2.3   6.3 $\pm$ 1.4 38.3 $\pm$ 2.9
TokenChange 25.4 $\pm$ 2.5 53.1 $\pm$ 2.9 25.4 $\pm$ 2.1 43.4 $\pm$ 2.3   6.1 $\pm$ 1.4 36.5 $\pm$ 2.8
EnsembleR (MA+VP) 62.1 $\pm$ 2.8 45.8 $\pm$ 2.9 59.6 $\pm$ 2.3 47.2 $\pm$ 2.4 51.2 $\pm$ 2.9 38.1 $\pm$ 2.9
EnsembleR (MA+TC) 65.8 $\pm$ 2.8 48.9 $\pm$ 2.9 58.8 $\pm$ 2.3 47.2 $\pm$ 2.4 51.1 $\pm$ 2.9 40.3 $\pm$ 2.9
EnsembleR (VP+TC) 22.6 $\pm$ 2.4 50.7 $\pm$ 2.9 29.2 $\pm$ 2.1 44.2 $\pm$ 2.4   7.1 $\pm$ 1.5 40.9 $\pm$ 2.9
EnsembleR (All) 62.7 $\pm$ 2.8 51.6 $\pm$ 2.9 60.4 $\pm$ 2.3 47.2 $\pm$ 2.4 50.2 $\pm$ 2.9 37.1 $\pm$ 2.8
EnsembleC (All) 39.1 $\pm$ 2.8 55.5 $\pm$ 2.9 42.4 $\pm$ 2.3 46.9 $\pm$ 2.4 24.4 $\pm$ 2.5 37.2 $\pm$ 2.8
Table 7: Input- and output-based evaluation results of the methods and their ensembles, over different feature types and models, averaged across model layers, along with their respective 95% confidence intervals. For GPT-2 small SAE features we take ones with width 32k. We denote MA for MaxAct, VP for VocabProj, TC for TokenChange, and EnsembleR and EnsembleC for the raw and concatenation based ensembles.
@@ Appendix C Additional Evaluation Results
Appendix D Computational Cost Analysis
The computational cost of each method is a key factor to consider when selecting a method for generating descriptions. In our analysis, we computed the FLOPs required by each method to generate a description for every single MLP feature in Gemma-2 2B (results in Table 4). When calculating the FLOPs required for a single forward pass, we rely on the heuristic FLOPs $\approx$ 6N plus embedding FLOPs, where N is the total number of non-embedding model parameters Kaplan et al. (2020); Anil et al. (2023). The results show that even when using a small sample for MaxAct—25k sequences of 128 tokens each, as used by Neuronpedia—alternative methods are 2-3 orders of magnitude more compute-efficient. When using larger samples that more accurately represent a model’s training data, such as The Pile Gao et al. (2020), the difference reaches 7-8 orders of magnitude. Lastly, computational cost for analysing SAE features results in an increase of roughly one order of magnitutde across the board, while maintaining the same relative differences between methods.
Appendix E Human Evaluations
To lend credence to our use of an LLM-judge and assess how well LLM-generated feature descriptions align with human judgment, we conducted two human evaluations.
E.1 Justifying Using an LLM as a Judge
To justify our use of an LLM-as-a-judge in the output-based evaluation, we apply the alternative annotator test proposed by Calderon et al. (2025). Following their procedure, we use three human annotators (graduate students) and a set of 100 randomly selected feature examples, evenly split between Llama-3.1 8B and Gemma-2 2B. For each feature, human annotators were given feature descriptions generated by VocabProj, and the three text sets $\mathcal{T}_{\mathbf{v}_{f}}$, $\mathcal{T}_{\mathbf{v}_{f}^{\prime}}$, and $\mathcal{T}_{\mathbf{v}_{f}^{\prime\prime}}$. Each annotator then indicated which of the three sets matches the given description, as per the output-based metric. Consistent with Calderon et al. (2025), we set $\epsilon=0.15$ to reflect our use of graduate student annotators. The analysis yielded a winning rate $\omega=1$ with p-value 0.03, supporting our use of an LLM-as-a-judge.
E.2 Evaluating LLM Generated Descriptions
To evaluate how well LLM-generated feature descriptions align with human judgment, we tasked human annotators (6 graduate students) with scoring their faithfulness with respect to (a) input-faithfulness: what activates the feature and (b) output-faithfulness: how the feature affects the model’s outputs. The instructions provided to the annotators are shown in Figure 6. We collected annotations for feature descriptions generated by MaxAct, VocabProj, and Ensemble Raw (All) for 100 randomly selected SAE features from Gemma-2 2B.
Figure 5 shows the results, where the overall trends align well with our proposed input- and output-based evaluations, discussed in §5.2. MaxAct performs better on the input evaluation, VocabProj on the output evaluation, and Ensemble Raw (All) performs best on both. However, VocabProj performed slightly worse than expected on the output evaluation. This discrepancy may stem from the difficulty humans face in evaluating a feature’s effect on text generation, as it requires detecting subtle changes across multiple texts. Indeed, in the annotator test conducted in §E, the judge LLM outperformed human annotators, supporting this claim. Furthermore, MaxAct’s success in the input evaluation could be influenced by the descriptions being derived from the same data used for comparison, potentially biasing results in its favor. Nonetheless, these findings reinforce the claims in §5.2, that input-centric methods perform better on input-based evaluations, output-centric methods on output-based ones, and ensembles perform best on both.
Appendix F Additional Details and Examples for Dead Feature Analysis
F.1 Generating Candidate Prompts
@@
F.2 Dead Feature Revival Example
@@
Appendix G Additional Examples for Qualitative Analysis
@@
Appendix H Resources and Packages
In our experiments, we used models, data and code from the following packages: transformers Wolf (2019), datasets Lhoest et al. (2021), TransformerLens Nanda and Bloom (2022) and SAELens Joseph Bloom and Chanin (2024). The authors also made use of AI models, specifically ChatGPT, for implementing specific helper functions. All of the experiments were conducted using a single A100 80GB or H100 80GB GPU.
@@ Appendix H Resources and Packages
Figure 14: An example of a steered text set for the output-based metric.
@@
Figure 16: Higher clamping value when steering feature with description “purchasing activities, including buying, viewing, and downloading products”, leading to degenerate text.